# Optimizing Frequent Checkpointing via Low-Cost Differential for Distributed Training Systems

### v2

@@ VIII Evaluation
  • How fast do LowDiff train under high checkpointing frequency compared to other methods? (Exp. 1)
  • How fast do LowDiff+ train under non-compression scenarios? (Exp. 2)
  • Can LowDiff and LowDiff+ achieve a shorter wasted time? (Exp. 3)
  • Does LowDiff and LowDiff+ accelerate the recovery process? (Exp. 5)
  • Can LowDiff reduce the storage overhead? (Exp. 7)
@@ VIII-A Experimental Setup
Baselines. We evaluate two state-of-the-art checkpointing strategies: CheckFreq [mohan2021checkfreq] and Gemini [wang2023gemini], as described in §II-B. The experiments follow the default configurations according to their original papers. Additionally, we assess the performance of the Naïve DC method, which computes the differences in model states between iterations and checkpoints these differentials, as Check-N-Run [eisenman2022check] does in recommendation systems.
@@ VIII-B Performance Evaluation
Figure 11: (Exp. 1) Training time of different models with different checkpointing strategies.
Experiment 1 (Training time of LowDiff). We evaluate the training time of LowDiff under different checkpointing strategies on the previously introduced DNN models, with the checkpointing frequency set to once per iteration. The experiment is conducted in the presence of gradient compression. For each model, we run 1,000 training iterations on A100 servers and record the total training time. The W/O CKPT setting represents training without checkpointing, which serves as the upper bound of achievable training speed.
Figure 11 shows that the training time of LowDiff on all seven training tasks is close to that of W/O CKPT and significantly smaller than those of other solutions. Compared to W/O CKPT, LowDiff’s training time increases by only 2.4%-3.1%, while the other checkpointing methods’ training time increases by 8.1%-891%.
Specifically, we elaborate on LowDiff’s benefits over the other methods and analyze the reasons behind these benefits:
@@
Figure 12: (Exp. 2) Training time without compression.
Experiment 2 (Training time of LowDiff+). To isolate the impact of gradient compression, we further evaluate the training time of LowDiff+ in a setting without gradient compression. All other experimental configurations, including the checkpointing frequency (per iteration), the selected DNN models, the number of training iterations (1,000), and the hardware environment (A100 servers), remain identical to Exp. 1.
Figure 12 shows that the training time of LowDiff+ across all tasks exhibits a slight increase compared to the W/O CKPT baseline, yet it remains significantly lower than that of other checkpointing solutions. Specifically, LowDiff+’s training time increases by 7.2%–9.1% relative to W/O CKPT. As expected, this overhead is higher than the 2.4%-3.1% observed for LowDiff in Exp. 1. This increase represents the necessary and expected performance trade-off for handling uncompressed, high-precision gradients, which was the primary motivation for the LowDiff+ design. The overhead is mainly due to frequent and large-volume gradient transfers between GPU and CPU memory, occupying PCIe bandwidth.
Despite this minor overhead, in the high-frequency checkpointing scenario without gradient compression, LowDiff+ consistently achieves the lowest training time among all checkpointing methods. For the GPT2-L training task, it reduces the training time by 51.8% compared to Gemini and by 81.7% compared to CheckFreq, demonstrating its superior efficiency in settings where frequent non-compression checkpoints are required.
Experiment 3 (Wasted time). We evaluate the wasted time of various checkpointing methods in the event of failures. Failures were simulated during training by adhering to a fixed Mean Time Between Failures (MTBF) metric. Using three MTBF values at 0.5, 1, and 2 [gupta2024just], respectively, we evaluate the wasted time of Naïve DC, CheckFreq, Gemini, and LowDiff when training GPT2-S. Note that LowDiff’s checkpointing optimal configuration is obtained based on Equation (5) in §V-C. For LowDiff+, we evaluate its wasted time under both software and hardware failures, where the corresponding checkpoints can be accessed throught snapshot and persist ( LowDiff+ (S) and LowDiff+ (P), respectively).
@@
For LowDiff+, the wasted time due to software failures is 3.7%–5.1% lower than that of LowDiff. This improvement comes from LowDiff+ ’s in-memory checkpointing, which enables fast recovery using the full model state stored in CPU memory. The detailed recovery times of LowDiff+ (S) are presented in Exp. 5. In contrast, the wasted time from hardware failures is slightly higher than that of LowDiff, due to LowDiff+ ’s lower checkpointing frequency, but still lower than that of CheckFreq and Gemini. Details of the checkpointing frequency for LowDiff+ (S) and LowDiff+ (H) are presented in Exp. 6.
Figure 13: (Exp. 3) Wasted time under different MTBFs.
@@
LowDiff+ (S) also performs per-iteration in-memory checkpointing, enabled by its efficient layer-wise gradient reuse and snapshotting approach. LowDiff+ (P) achieves per-iteration persistence frequency in smaller models such as ResNet-101, but increases to 3 iterations for larger models like GPT2-L due to the limited bandwidth of SSDs. Nevertheless, its checkpointing intervals remain the shortest compared to Naïve DC, CheckFreq, and Gemini.
@@
Experiment 5 (Recovery time). We evaluate the recovery time of different recovery strategies under different full checkpointing frequency in GPT2-S. We compare traditional checkpointing without differential checkpoint (Baseline) (i.e., Torch.save [paszke2019pytorch]), Naïve DC, LowDiff with parallel recovery (see §VII), and LowDiff+ from software failures (LowDiff+ (S) for short). Note that recovery from hardware failures in LowDiff+ is the same as in the baseline, since both reload the model from storage and resume training from the last persisted checkpoint, as determined by the checkpointing frequency.
[…]
@@
As shown in Figure 17, LowDiff enables per-iteration checkpointing for GPT2-S training across the entire range of $\rho\in[0.001,0.1]$ due to its smaller enough gradient size for checkpointing to be overlapped with training within one iteration. For GPT2-L, LowDiff supports per-iteration checkpointing when $\rho\in[0.001,0.075]$, while at $\rho=0.1$, the checkpoint frequency increases to 2 iterations because the increasing gradient size makes the checkpointing take two iterations to be overlapped with training. The result shows LowDiff ensures frequent checkpointing (intervals ¡ 3 iterations) under the common compression ratio.
@@ VIII-C System Scalability
Experiment 9 (Scaling to frequent failures). We conduct experiments on the performance of LowDiff and LowDiff+ in extreme scenarios characterized by more frequent failures of the training process with MTBF ranging from 0.1 to 5 hours. Using the effective training time ratio metric as defined in Gemini [wang2023gemini], which measures the percentage of time spent on productive training progress within a given period of time, we evaluate the effective training time ratios of Torch.save (baseline), CheckFreq, Gemini, LowDiff, and LowDiff+. Note that LowDiff+ has two recovery modes: LowDiff+ (S) for software failures (recovering from CPU memory) and LowDiff+ (P) for hardware failures (recovering from persistent storage).
Figure 18 demonstrates that LowDiff+ (S) consistently achieves the highest effective training time ratio across all scenarios, especially under environments with frequent failures. This is because its near-instantaneous recovery from in-memory checkpoints (as shown in Exp. 5) almost completely eliminates recovery overhead, compensating for its slightly higher steady-state overhead compared to LowDiff. LowDiff achieves the second-highest ratio. LowDiff+ (P) also consistently outperforms the other baselines, though its reliance on slower disk-based recovery places it slightly below LowDiff. For example, when the MTBF is set to 0.3 hours, LowDiff+ (S) leads with an effective training time ratio of 94.0%, followed closely by LowDiff at 92%. LowDiff+ (P) achieves 86.8%, which is still significantly higher than Gemini (81%) and Checkfreq (75.9%). These results highlight the robustness of our approaches, with LowDiff+ (S) offering the best performance for common software failures.
@@
Figure 19 shows that, as the number of GPUs increases, the effective training time ratio decreases; however, LowDiff still maintains a 98% proportion of the effective training time ratio, and LowDiff+ achieves 96%, while the metric for the other method declines rapidly and can only reach around 90%. This highlights the strong potential of LowDiff and LowDiff+ in large-scale training.
@@ IX Conclusion
LowDiff is an efficient frequent checkpointing framework that addresses the challenges of the high costs of differential checkpointing by reusing the compressed gradients generated during training. LowDiff also designs a batched gradient writing optimization scheme and an optimal configuration tuning scheme. To enhance the gradient reuse strategy in non-compression scenarios, we design LowDiff+. LowDiff+ uses a layer-wise gradient reusing and snapshotting strategy and a CPU-based synchronous persistence strategy, thus enabling checkpointing to decouple from training. Our experiments on a variety of models and GPU generations demonstrate the efficiency of LowDiff and LowDiff+ in terms of the training time, wasted time, checkpointing time, and recovery time under the high-frequency checkpointing scenarios.

### v3

@@ VIII Evaluation
  • How fast do LowDiff and LowDiff+ train under high checkpointing frequency with gradient compression? (Exp. 1)
  • How fast do LowDiff and LowDiff+ train without gradient compression? (Exp. 2)
  • Can LowDiff and LowDiff+ achieve a shorter wasted time? (Exp. 3)
  • Does LowDiff and LowDiff+ accelerate the recovery process? (Exp. 5)
  • Can LowDiff reduce the storage overhead? (Exp. 7)
@@ VIII-A Experimental Setup
Comparative Methods. We evaluate two state-of-the-art checkpointing strategies: CheckFreq [mohan2021checkfreq] and Gemini [wang2023gemini], as described in §II-B. The experiments follow the default configurations according to their original papers. Additionally, we assess the performance of the Naïve DC method, which computes the differences in model states between iterations and checkpoints these differentials, as Check-N-Run [eisenman2022check] does in recommendation systems. For LowDiff+, we denote the snapshotting phase as LowDiff+(S) and the persistence phase as LowDiff+(P).
@@ VIII-B Performance Evaluation
Figure 12: (Exp. 1) Training time with gradient compression.
Experiment 1 (Training time with gradient compression). We evaluate the training time with gradient compression of LowDiff and LowDiff+ under different checkpointing strategies on the previously introduced DNN models, with the checkpointing frequency set to once per iteration. The experiment is conducted in the presence of gradient compression. For each model, we run 1,000 training iterations on A100 servers and record the total training time. The W/O CKPT setting represents training without checkpointing, which serves as the upper bound of achievable training speed.
Figure 12 shows that the training time of LowDiff and LowDiff+ on all seven training tasks is close to that of W/O CKPT and significantly shorter than that of other solutions. Compared to W/O CKPT, LowDiff and LowDiff+’s training time increases by only 2.4%-3.1% and 0.7%-2.9%, respectively, while the other checkpointing methods’ training time increases by 8.1%-891%.
Specifically, we elaborate on LowDiff and LowDiff+’s benefits over the other methods and analyze the reasons behind these benefits:
@@
  • For LowDiff+, its training time across all models is even shorter than LowDiff. Specifically, LowDiff+ reduces the training time by 0.8%, 1.5%, 1.7%, 1.6%, 1.1%, and 3.3% in ResNet-101, VGG-19, BERT-B, BERT-L, GPT2-S, and GPT2-L, respectively. While the compressed gradients are already small enough to be handled efficiently by LowDiff’s DC, limiting the gains in snapshotting and persistence of LowDiff+, the advantages of LowDiff+ become more pronounced in non-compression scenarios, as further discussed in Exp. 2.
Figure 13: (Exp. 2) Training time without gradient compression.
Experiment 2 (Training time without gradient compression.) We further evaluate the training time of LowDiff and LowDiff+ in a scenario where gradient compression is disabled, while keeping other settings the same as Exp. 1.
As shown in Figure 13, LowDiff+ consistently achieves the lowest training time among all evaluated checkpointing methods. Specifically, LowDiff+ consistently achieves the lowest training time among all checkpointing methods. For the GPT2-L training task, it reduces the training time by 42.4% compared to Gemini and by 81.2% compared to CheckFreq, demonstrating its efficiency in settings where frequent non-compression checkpoints are required.
Furthermore, the results highlight the effectiveness of the proposed enhancements in LowDiff+. In this non-compression setting, the original LowDiff incurs significant overhead, with training time increasing by 16.1%–28.1% compared to the W/O CKPT baseline across different models. In contrast, LowDiff+ limits this overhead to an average of only 5.1%-7.2%. Although this is higher than the 1.8% overhead observed in Exp. 1 (due to the increased bandwidth pressure from uncompressed gradients), the substantial improvement over LowDiff confirms that our layer-wise-resue snapshotting and incremental-merging persistence successfully mitigate the snapshotting and persistence issues in §VI-A.
Figure 14: (Exp. 3) Wasted time under different MTBFs.
Experiment 3 (Wasted time). We evaluate the wasted time of various checkpointing methods in the event of failures. Failures were simulated during training by adhering to a fixed Mean Time Between Failures (MTBF) metric. Using three MTBF values at 0.5, 1, and 2 [gupta2024just], respectively, we evaluate the wasted time of Naïve DC, CheckFreq, Gemini, and LowDiff when training GPT2-S. Note that LowDiff’s checkpointing optimal configuration is obtained based on Equation (5) in §V-C. For LowDiff+, we evaluate its wasted time under both software and hardware failures, where the corresponding checkpoints can be accessed through LowDiff+(S) and LowDiff+(P).
@@
[…]
@@
LowDiff+(S) also performs per-iteration in-memory checkpointing, enabled by its efficient layer-wise-reuse snapshotting approach. LowDiff+ (P) achieves per-iteration persistence frequency in smaller models such as ResNet-101, but increases to 3 iterations for larger models like GPT2-L due to the limited bandwidth of SSDs. Nevertheless, its checkpointing intervals remain shorter than those of Naïve DC and CheckFreq, and are nearly the same as those of Gemini.
@@
Experiment 5 (Recovery time). We evaluate the recovery time of different recovery strategies under different full checkpointing frequencies in GPT2-S. We compare traditional checkpointing without differential checkpoint as Baseline (i.e., Torch.save [paszke2019pytorch]), Naïve DC, LowDiff with parallel recovery (see §VII), and LowDiff+ from software failures (LowDiff+(S) for short). Note that recovery from hardware failures in LowDiff+ is the same as in the baseline, since both reload the model from storage and resume training from the last persisted checkpoint, as determined by the checkpointing frequency.
[…]
@@
As shown in Figure 18, LowDiff enables per-iteration checkpointing for GPT2-S training across the entire range of $\rho\in[0.001,0.1]$ due to its smaller enough gradient size for checkpointing to be overlapped with training within one iteration. For GPT2-L, LowDiff supports per-iteration checkpointing when $\rho\in[0.001,0.075]$, while at $\rho=0.1$, the checkpoint frequency increases to 2 iterations because the increasing gradient size makes the checkpointing take two iterations to be overlapped with training. The result shows LowDiff ensures frequent checkpointing (intervals <3 iterations) under the common compression ratio.
@@ VIII-C System Scalability
Experiment 9 (Scaling to frequent failures). We conduct experiments on the performance of LowDiff and LowDiff+ in extreme scenarios characterized by more frequent failures of the training process with MTBF ranging from 0.1 to 5 hours. Using the effective training time ratio metric as defined in Gemini [wang2023gemini], which measures the percentage of time spent on productive training progress within a given period of time, we evaluate the effective training time ratios of Torch.save (baseline), CheckFreq, Gemini, LowDiff, and LowDiff+.
Figure 19 demonstrates that LowDiff+ consistently achieves the highest effective training time ratio across all scenarios, especially under environments with frequent failures. This is because its fast recovery from in-memory full model state (as shown in Exp. 5) almost eliminates recovery overhead, compensating for its slightly higher steady-state overhead compared to LowDiff. LowDiff achieves the second-highest ratio. For example, when the MTBF is set to 0.3 hours, LowDiff+ leads with an effective training time ratio of 93.3%, followed closely by LowDiff at 92.1%. These results highlight the robustness of our approaches, with LowDiff+ offering the best performance for common software failures.
Figure 20 shows that, as the number of GPUs increases, the effective training time ratio decreases; however, LowDiff and LowDiff+ still maintain a nearly 98% proportion of the effective training time ratio, while the metric for the other method declines rapidly and can only reach around 90%. This highlights the strong potential of LowDiff and LowDiff+ in large-scale training.
@@ IX Conclusion
LowDiff is an efficient frequent checkpointing framework that addresses the challenges of the high costs of differential checkpointing by reusing the compressed gradients generated during training. LowDiff also designs a batched gradient writing optimization scheme and an optimal configuration tuning scheme. To enhance the gradient reuse strategy in non-compression scenarios, we design LowDiff+. LowDiff+ uses a layer-wise-reuse snapshotting strategy and an incremental-merging persistence strategy, thus enabling checkpointing to decouple from training. Our experiments on a variety of models and GPU generations demonstrate the efficiency of LowDiff and LowDiff+ in terms of the training time, wasted time, checkpointing time, and recovery time under the high-frequency checkpointing scenarios.